An ideal heat-engine is to be used in an environment where the ambient temperature is 31.5 °C. What should be the minimum temperature of the hot heat reservoir in order to reach at least 36.9 percent efficiency with the heat-engine? (Give your answer in Celsius.)
Answer #1


efficiency n = W/Q = (Qhot - Qcold)/Qhot


Q is proportional to temperatre T


percent efficiency = (Thot - Tcold)/Thot


given n = 36.9% = 0.369


Tcold = 31.5 + 273 = 304.5 K

0.369 = (Thot-304.5)/Thot

Thot = 482.57 K

Thot = 482.57-273 = 209.5 0C